### v4

# Optimal Estimation of Temperature

Abstract
Temperature of a finite-sized system fluctuates due to the thermal fluctuations. However, a systematic mathematical framework for measuring or estimating the temperature is still underdeveloped. Here, we incorporate the estimation theory in statistical inference to estimate the temperature of a finite-sized system and propose optimal estimation based on the uniform minimum variance unbiased estimation. Treating the finite-sized system as a thermometer measuring the temperature of a heat reservoir, we demonstrate that different optimal estimation of parameters yield different formulas of entropy, e.g., optimal estimation of inverse temperature (or temperature) aligns with the Boltzmann entropy (or Gibbs entropy). The optimal estimation leads to a achievable energy-temperature uncertainty relation and exhibits sample-size dependence, coinciding with their counterparts in nanothermodynamics. The achievable bound and the non-Gaussian distribution of temperature enable experimental testing in finite-sized systems.
I Introduction
Temperature fluctuations become increasingly apparent as the system size decreases. The significant fluctuations of temperature in finite-sized systems challenge the zeroth law of thermodynamics and the notion of a well-defined temperature for the finite-sized system. Such challenges have sparked a wide range of subsequent studies, including the definition of entropy and temperature in isolated systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], the existence of negative absolute temperatures [11, 12, 13, 14, 15, 16, 17, 18, 19, 20], high resolution thermometer (HRT) [21, 22, 23, 24], quantum calorimeter [25] or quantum thermometry [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], the energy–temperature uncertainty relation (ETU) [38, 39], the thermodynamics of finite-sized systems [40, 41, 42, 43, 44, 45], and nonequilibrium processes and stochastic thermodynamics [46, 47, 48, 49]. Despite considerable progress, conclusions drawn from specific viewpoints or scenarios are sometimes inconsistent, and a unified understanding of thermal fluctuations in finite-sized systems remains incomplete. This situation reflects the absence of a general mathematical framework for studying such fluctuations, which motivates our work to develop a more systematic approach.
Historically, Einstein firstly investigated equilibrium fluctuations by relating the probability of macroscopic states to entropy through an inversion of the Boltzmann’s entropy formula [50, 51, 52, 53]. Bohr suggested that there should exist a form of complementarity between temperature and energy in thermodynamics similar to that of position and momentum in quantum theory [38, 39]. Then, Mandelbrot linked the ETU ($\beta$ denotes the inverse temperature of the system and $E$ denotes the energy of the system)
$\Delta\beta\Delta E\geq 1$ (1)
to the Cramér–Rao bound in estimation theory through the polemic exchange with Feshbach and Kittel [54, 55, 56, 57], where $\Delta(\cdot)=\sqrt{\langle\cdot^{2}\rangle-\langle\cdot\rangle^{2}}$. This practice originated from 1956 [58] when Mandelbrot applied the estimation theory to the derivation of statistical thermodynamics from purely phenomenological principles (also see [59, 60, 61, 62]). Experimentally, Gaussian-type temperature fluctuations have been observed in paramagnetic salt HRT [21, 22, 23], ferromagnetic alloy HRT [24], and a nanoscale electron calorimeter [25]. They verified that the temperature of the system exhibits fluctuations, while the temperature of the reservoir remains fixed.
In this article, aiming to extend the study beyond the Gaussian-type fluctuation near the thermodynamic limit, we develop a general mathematical framework for thermal fluctuations based on estimation theory. By proposing the concept, optimal estimation of temperature, we rigorously describe temperature fluctuations in a finite-sized system within the framework of statistical inference. This approach provides a systematic perspective for studying the thermodynamics of finite-sized systems.
In estimation theory, unbiasedness and efficiency are the two important criteria for evaluating estimators. An estimator $\hat{\theta}$ of parameter $\theta$ is called “unbiased”, if its mean value satisfies [63]
$\langle\hat{\theta}\rangle=\theta,\quad\text{for}\quad\theta\in\Theta,$ (2)
where $\Theta$ is the range of all possible parameter values. If the mean squared errors of two estimators $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$ satisfy
$\langle(\hat{\theta}_{1}-\theta)^{2}\rangle\leq\langle(\hat{\theta}_{2}-\theta)^{2}\rangle,\quad\text{for}\quad\theta\in\Theta,$ (3)
then $\hat{\theta}_{1}$ is called to be more efficient than $\hat{\theta}_{2}$ [63]. In experimental terms, unbiasedness corresponds to accuracy (the absence of systematic error), while efficiency corresponds to precision (the minimization of statistical uncertainty). Therefore, an optimal estimator is the most efficient estimator among all unbiased estimators for any $\theta\in\Theta$, i.e., the uniform minimum variance unbiased estimation (UMVUE).
For systems without upper energy bound, we find a one-to-one link between Boltzmann (Gibbs) entropy $S_{\text{B}}=\ln[\sigma(E)\epsilon]$ $(S_{\text{G}}=\ln\Omega(E))$ and the optimal estimation of inverse temperature (temperature), i.e., Boltzmann temperature $\hat{\beta}_{B}\equiv\partial S_{B}/\partial E$ (Gibbs temperature $\hat{T}_{G}\equiv(\partial S_{G}/\partial E)^{-1}$). Here, $\sigma(E)$ is the density of the states of the system at energy $E$, $\epsilon$ is the width of a narrow energy interval, and $\Omega(E)=\int_{E_{\text{min}}}^{E}\text{d}E^{{}^{\prime}}\sigma(E^{{}^{\prime}})$ is the volume in phase space enclosed by a hypersurface of constant energy $E$ [64], where $E_{\text{min}}$ denotes the lower bound of system’s energy and we set Boltzmann constant $k_{B}=1$. Hence, different formulas of entropy apply to optimal estimation of different parameters, rather than being contradictory. For systems with upper energy bound, the Boltzmann temperature still yields optimal estimation regardless of whether the temperature is positive or negative, while the Gibbs temperature yields optimal estimation only for positive temperatures.
The variance of the optimal estimator of temperature is the infimum among the variances of all unbiased estimators of temperature, thus yielding an achievable ETU. In the large-$N$ limit, it leads to a refined ETU
$\Delta\beta\Delta E\geq\Delta\hat{\beta}_{B}\Delta E=1+\frac{1}{4}(\mu_{3})^{2},$ (4)
where $\mu_{3}=\langle(E-\left\langle E\right\rangle)^{3}\rangle/(\Delta E)^{3}$ denotes the skewness of the canonical distribution, and $\hat{\beta}_{B}$ is shown to be the optimal estimator of the inverse temperature in the following paper.
In repeated sampling, the optimal estimation of temperature exhibits sample-size dependence, coinciding with their counterparts in nanothermodynamics [40, 41, 42, 43, 44, 45]. Actually, the “sampling ensemble” can be regarded as the statistical interpretation of the replica trick in Hill’s nanothermodynamics. And a Gaussian-type temperature fluctuation naturally emerges from a general temperature distribution due to the central limit theorem (CLT).
Compared with quantum thermometry, which mainly focuses on measuring temperature via quantum coherence in few-body systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], our work provides a framework for describing temperature fluctuations in finite-sized systems, offering an alternative approach to achieve more accurate temperature estimates in many-body systems.
II Estimation of temperature in single sampling
Developing Mandelbrot’s viewpoint about temperature fluctuations [56, 58, 59, 60], i.e., “temperature for systems-in-isolation should be viewed as a statistical estimate of the parameter of a conjectural canonical distribution”, we treat a finite-sized system in canonical ensemble as a thermometer of a reservoir. The $N$-particle system, with $i\in D$ denoting its microstate, contacts a real or hypothetical heat reservoir with temperature $T$, where $D$ denotes the set of energy level indices. At equilibrium, the probability of the system at state $i$ obeys the Boltzmann-Gibbs formula:
$P_{i}=\frac{d_{i}}{Z}e^{-\beta E_{i}},$ (5)
where $\beta=T^{-1}$ denotes the inverse temperature, $E_{i}$ $(d_{i})$ denotes the energy (degeneracy) of state $i$ and $Z=\sum_{i\in D}d_{i}e^{-\beta E_{i}}$ is the partition function.
Let $E_{\text{max}}$ denote the upper bound of system’s energy, $\hat{\beta}^{(k)}$ ($\hat{T}^{(k)}$) denote the unbiased estimator of $\beta^{k}$ ($T^{k}$), where $k$ is an integer. We discuss the unbiased estimation of temperature in two cases.
Firstly, when $E_{\text{max}}=\infty$, one option of $\hat{\beta}^{(k)}$ reads (see Appendix A):
$\hat{\beta}^{(k)}=\begin{cases}\frac{\text{d}^{k}_{E_{i}}\sigma(E_{i})}{\sigma(E_{i})}&\text{if }\sigma(E_{i})\neq 0\\ 0&\text{if }\sigma(E_{i})=0\end{cases},$ (6)
where $\sigma(E)\equiv\sum_{i\in D}\delta(E-E_{i})d_{i}$ denotes the density of states , $\text{d}_{E}^{0}\sigma(E)\equiv\sigma(E)$ and [65]
$\text{d}^{k}_{E}\sigma(E)\equiv\begin{cases}\frac{\text{d}^{k}\sigma(E)}{\text{d}E^{k}}&\text{for }k\textgreater 0\\ \int_{E_{\text{min}}}^{E}\text{d}E^{{}^{\prime}}\text{d}_{E^{\prime}}^{k+1}\sigma(E^{\prime})&\text{for }k\textless 0\end{cases}.$ (7)
In the derivation, we have assumed the density of states $\sigma(E)$ satisfies
$\left.\text{d}^{l}_{E}\sigma(E)\right|_{E=E_{\text{min}}}=0$ (8)
for integer $l\in[0,k)$, which holds in classical systems but not in quantum systems. This is relevant to the third law of thermodynamics [66].
When the above assumption (8) is invalid, the unbiased estimator of $\beta^{k}$ for $k<0$, as well as that of $\beta$ (except for an exponentially decaying bias with $N$), are still given in Eq. (6). In FIG.1, we illustrate this bias in an analogous low-temperature Fermi gas system using relative bias $\Phi_{\beta}=|(\langle\hat{\beta}^{(1)}\rangle-\beta)/\beta|$.
Figure 1: Relative bias of $\beta$ varies as NN when $E_{\text{max}}=\infty$. Here, $\sigma(E)=e^{\sqrt{N(E-E_{\text{min}})}}$ is analogous to the density of states of the low-temperature ideal fermi gas (see Appendix B), where the value of $E_{\text{min}}$ does not affect the results.
It is worth noting that $k$ has an upper bound (in this paper we mainly consider the case $k=\pm 1$), depending on the degrees of freedom of the system. As a example, taking the density of states $\sigma(E)=E^{3N/2-1}$ of a classical ideal gas as an example [67], $\hat{\beta}^{(k)}$ in Eq. (6) is
$\hat{\beta}^{(k)}=\frac{\Gamma\left(\frac{3N}{2}\right)}{\Gamma\left(\frac{3N}{2}-k\right)}E^{-k},$ (9)
where $k\leq 3N/2-2$ and $\Gamma$ denotes the Gamma function [68].
When considering the density of states of a quantum many-body system, an issue arises of how to define the derivatives and integrals in Eqs. (6) and (7) since the density of states is discrete. In practice, this is resolved by approximating the discrete density of states with a continuous function, as commonly done in statistical mechanics. For instance, in Section 1.4 of the textbook [69], the author carefully discusses how to approximate the density of states and the phase volume of a quantum ideal gas using a continuous function, as well as the effect of boundary conditions on the approximation.
Secondly, when $E_{\text{max}}$ is finite, the unbiased estimator of inverse temperature is still given in Eq.(6), regardless of whether the inverse temperature is positive or negative. However, it exhibits an exponentially decaying bias with increasing $N$. In FIG.2, we illustrate this bias in an $N$ non-interacting two-level systems, showing that it leads to a relative bias $\Phi_{\beta}\sim\text{e}^{-\alpha_{1}N}$, where the decay rate $\alpha_{1}=\ln(1+\text{e}^{-|\beta|\varepsilon})$ (see Appendix C), and $\varepsilon$ denotes the spacing of energy levels.
Figure 2: Relative bias of $\beta$ varies as NN when $E_{\text{max}}$ is finite. Here, $\sigma(E)$ is the density of states of an $N$ non-interacting two-level system (see Appendix C) and we set the spacing of energy levels $\varepsilon=1$.
For positive temperatures, the unbiased estimator of temperature remains given by Eq. (6), except for a similar exponentially decaying bias. In contrast, for negative temperatures, using Eq. (6) to estimate the system’s temperature results in a bias that grows exponentially, which demonstrates that Gibbs temperature cannot be used to estimate the temperature of a negative-temperature system. In FIG.3, we illustrate these biases in an $N$ non-interacting two-level systems, showing that they lead to a relative bias $\Phi_{T}=|(\langle\hat{\beta}^{(-1)}\rangle-T)/T|\sim\text{e}^{-\alpha_{2}N}$, where the decay (increase) rate $\alpha_{2}=\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2})$ (see Appendix C). The above indicates that an optimal estimator for the temperature of a negative-temperature system likely does not exist. Moreover, we extend the unbiased estimator of $\beta^{\alpha}$ to $\alpha\in\mathbb{R}$ (see Appendix D).
Figure 3: Relative bias of $T$ varies as NN when $E_{\text{max}}$ is finite. Here, the density of states is the same as FIG.2 (see Appendix C). (a) $T>0$. (b) $T<0$.
In the following, we prove one of the main results of this paper: the unbiased estimators in Eq. (6) are optimal, i.e., the UMVUEs of temperature in canonical ensemble. First, we introduce an essential theorem in statistical inference, the Rao-Blackwell-Lehmann-Scheffé theorem (RBLS) [63, 70, 71]:
Theorem 1.
If $\hat{\theta}$ is an unbiased estimator of $\theta$ and $T$ is a sufficient statistic for $\theta$, The expectation of $\hat{\theta}$ conditional on $T$, i.e., $\tilde{\theta}=\langle\hat{\theta}|T\rangle$ is
1. a vaild estimator for $\theta$ (not dependent on $\theta$)
2. unbiased
3. of lesser or equal variance than that of $\hat{\theta}$, for all $\theta$
Additionally, if the sufficient statistic is complete, then the unbiased estimator is the UMVUE.
A theorem [63, 70, 71] following the RBLS theorem is
Theorem 2.
Let $T$ be a complete sufficient statistic for a parameter $\theta$, and let $\tilde{\theta}(T)$ be any estimator based only on $T$. Then $\tilde{\theta}(T)$ is the unique UMVUE of its expected value.
Theorem 2 tells us that the UMVUE of temperature in the canonical ensemble is a function of its complete sufficient statistic. Meanwhile the Boltzmann–Gibbs formula in Eq. (5) belongs to the exponential family in statistical inference, and its complete sufficient statistic is $E_{i}$ [63]. Because the estimators $\hat{\beta}^{(k)}$ in Eq. (6) are not only unbiased but also a function of $E_{i}$, the estimators of $\hat{\beta}^{(k)}$ are the UMVUEs of $\beta^{k}$ according to Theorem 2. Moreover, no simple functional relationship exists among the UMVUEs of different moments of $\beta$, indicating that a unique definition of temperature does not exist in finite-sized systems.
Especially, for $k=\pm 1$, we have
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\langle\hat{\beta}_{B}\rangle=\beta.\\ \langle\hat{\beta}^{(-1)}\rangle&=\langle\hat{T}_{G}\rangle=T.\end{split}$ (10)
Hence, the Boltzmann entropy $S_{B}$ and the Gibbs entropy $S_{G}$ correspond to the UMVUE of inverse temperature and temperature respectively. It is worth mentioning that $\hat{\beta}_{B}$ and $\hat{T}_{G}$ have been discussed in previous studies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. The starting points of these studies are almost entirely based on the four laws of thermodynamics, which are established on empirical facts about systems in the thermodynamic limit. However, for finite-sized systems, whether these empirical facts and the four laws of thermodynamics are still valid is doubtful and unconfirmed. In contrast, we base our approach on the mathematical theorems in statistical inference, which is rigorous.
In our framework, the optimized scheme for obtaining the optimal estimation of temperature is as follows:
  1. 1.
    Determine the expressions for $\sigma$ and $\Omega$, which are either known before the measurement or can be obtained using the method in Section III. This method relates $\sigma$ to the partition function and its derivatives in the large-$N$ limit, allowing us to calculate $\sigma$ and $\Omega$ once the partition function is known.
  2. 2.
    Once the expressions for $\sigma$ and $\Omega$ are available, the expression for the estimator of temperature (Eqs. (6) and (7)) can be determined.
  3. 3.
    Measure the energy of the system and substitute it into Eqs. (6) and (7) to obtain the optimal estimation of the temperature.
The optimal estimate of temperature represents the temperature of the finite-sized system in this measurement. In practice, by substituting the measured energy into Eqs (6) and (7), the resulting estimator corresponds to the temperature of the finite-sized system for that particular measurement according to Mandelbrot’s viewpoint [56, 58, 59, 60]. This approach provides a theoretical foundation for analyzing temperature fluctuations. In fact, that is how temperature fluctuations are measured in experiments. For example, in the paramagnetic salt HRT [21, 22, 23], temperature is obtained by measuring changes in the magnetization. In our theoretical framework, this is not optimal and can be improved by measuring energy instead.
III Estimation of temperature for large-$N$ systems
The UMVUE of $T$ and $\beta$, $\hat{T}_{G}$ and $\hat{\beta}_{B}$, depend on the expression of $\sigma(E)$. For large-$N$ systems, we approximate the expression of $\sigma(E)$ through the partition function $Z$ and its derivatives, which can be obtained through measuring thermodynamic quantities. Using the Laplace approximation (saddle point approximation) [72, 73] or the Darwin–Fowler method in Sec. 9.1 of [67] (see Appendix E), we obtain
$\ln\sigma(\left\langle E\right\rangle)=S-\ln\sqrt{2\pi\left(-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)}+O(N^{-1}),$ (11)
where $S=\ln Z-\beta\partial\ln Z/\partial\beta$ is the thermodynamics entropy in a canonical ensemble.
Taking the derivative of Eq. (11) with respect to $E$ on both sides, we have
$\hat{\beta}_{B}(E)=\hat{\beta}_{L}(E)+\frac{1}{2}\frac{\text{d}^{2}_{E}\hat{\beta}_{L}(E)}{\text{d}_{E}\hat{\beta}_{L}(E)}+O(N^{-2}),$ (12)
where $\hat{\beta}_{L}(E)$ (given by $\left.\partial\ln Z/\partial\beta\right|_{\beta=\hat{\beta}_{L}(E)}=E$) is the the first-order moment (also the maximum likelihood) estimator of $\beta$.
Because $\hat{\beta}_{B}$ is the UMVUE of $\beta$, its variance is the infimum of among variances of unbiased estimators of $\beta$, i.e.,
$\Delta\hat{\beta}\Delta E\geq\Delta\hat{\beta}_{B}\Delta E\geq 1,$ (13)
where the variance of energy in the canonical ensemble $(\Delta E)^{2}=T^{2}C$ is also the Fisher information of parameter $\beta$ and $C$ denotes the heat capacity. Eq. (13) is the achievable ETU for finite $N$, which is tighter than the usual ETU.
In the large-$N$ limit, the optimal estimation of temperature leads to a refined ETU using Eq. (12) (see Appendix E):
$\Delta\hat{\beta}\Delta E\geq\Delta\hat{\beta}_{B}\Delta E=1+\frac{1}{4}(\mu_{3})^{2}+O(N^{-2}).$ (14)
In FIG. 4, the achievable ETU approach the refined ETU as $N$ increases. As $N\to\infty$, the skewness approaches 0 due to the CLT. The Gaussian-type temperature fluctuation and the usual ETU are recovered.
Figure 4: ETU varies as $1/N$. Here, $\sigma(E)=E^{\frac{3N}{2}-1}$ is the density of states of the classical ideal gas (see Appendix F). We set the inverse temperature $\beta=1$.
Similarly, $\hat{T}_{G}$ and $\hat{T}_{L}=\hat{\beta}_{L}(E)^{-1}$ satisfy
$\hat{T}_{G}(E)=\hat{T}_{L}(E)-\frac{\hat{T}_{L}(E)^{2}}{2}\frac{\text{d}^{2}_{E}\hat{T}_{L}(E)}{\text{d}_{E}\hat{T}_{L}(E)}+O(N^{-2}).$ (15)
Because $\hat{T}_{G}$ is the UMVUE of $T$, its variance is the infimum of among variances of unbiased estimators of $T$, i.e.,
$\frac{\Delta\hat{T}\Delta E}{T^{2}}\geq\frac{\Delta\hat{T}_{G}\Delta E}{T^{2}}\geq 1.$ (16)
In the large-$N$ limit, the optimal estimation of temperature leads to a refined $\Delta T\Delta E/T^{2}$ using Eq. (15) (see Appendix E):
$\frac{\Delta\hat{T}\Delta E}{T^{2}}\geq\frac{\Delta\hat{T}_{G}\Delta E}{T^{2}}=1+\frac{T^{2}}{4C^{3}}\left(\frac{\text{d}C}{\text{d}T}\right)^{2}+O(N^{-2}).$ (17)
Also, the second term on the right-hand side of Eq. (17) approaches 0 as $N\to\infty$ due to the CLT.
As an example, we consider classical ideal gas to illustrate the estimation of temperature (see Appendix F).
IV Estimation of temperature in repeated sampling
In single sampling (obtaining one outcome of the system energy), the finite-sized system exhibits significant fluctuations. To improve the precision of measurement, we discuss the UMVUE of temperature in repeated sampling with the sample size $M$. Let $\bm{i}$$\left(i_{1},i_{2},...,i_{M}\right)$ denote the microstates in repeated sampling. The $M$-independent sampling is equivalent to a single sampling of $M$-identical non-interacting systems. Here, we call the $M$-identical non-interacting systems “$M$-sampling ensemble”.
At equilibrium, the probability of the system at state $\bm{i}$ still obeys the Boltzmann-Gibbs formula:
$P_{\bm{i}}=\frac{d_{\bm{i}}}{Z^{M}}\text{e}^{-\beta E_{\bm{i}}},$ (18)
where $E_{\bm{i}}=\sum_{n=1}^{M}E_{i_{n}}\quad(d_{\bm{i}}=\prod_{n=1}^{M}d_{i_{n}})$ denotes the total energy (degeneracy) of state $\bm{i}$.
Similar to the single-sampling case, the unbiased estimator of $\beta^{k}$ with sample size $M$ is
$\hat{\beta}_{M}^{(k)}=\begin{cases}\frac{\text{d}^{k}_{E_{\bm{i}}}\sigma_{M}\left(E_{\bm{i}}\right)}{\sigma_{M}\left(E_{\bm{i}}\right)}&\text{if }\sigma_{M}\left(E_{\bm{i}}\right)\neq 0\\ 0&\text{if }\sigma_{M}\left(E_{\bm{i}}\right)=0\end{cases},$ (19)
where $\sigma_{M}(E)\equiv\sum_{\bm{i}}\delta\left(E-E_{\bm{i}}\right)d_{\bm{i}}$ denotes the density of states of the sampling ensemble. Because Eq. (18) also belongs to the exponential family, Eq. (19) is the UMVUE of $\beta^{k}$ in repeated sampling for system without upper energy bound, according to the RBLS theorem [63, 70, 71]. For system with upper energy bound, the result is similar to that obtained earlier. Especially, for $k=\pm 1$, we have
$\langle\hat{\beta}_{M}^{(1)}\rangle$ $=\langle\hat{\beta}_{B,M}\rangle=\left\langle\frac{\partial\ln(\sigma_{M}\epsilon)}{\partial E}\right\rangle=\beta,$ (20)
$\langle\hat{\beta}_{M}^{(-1)}\rangle$ $=\langle\hat{T}_{G,M}\rangle=\left\langle\left(\frac{\partial\ln\Omega_{M}}{\partial E}\right)^{-1}\right\rangle=T,$ (21)
which motivates us to generalize the definitions of Boltzmann entropy and Gibbs entropy in the “$M$-sampling ensemble”,
$S_{B,M}(E)$ $=\ln[\sigma_{M}(E)\epsilon],$ (22)
$S_{G,M}(E)$ $=\ln\Omega_{M}(E),$ (23)
where $\Omega_{M}(E)=\int_{E_{\text{min}}}^{E}\sigma_{M}(E^{\prime})\text{d}E^{\prime}$ is the volume in phase space enclosed by a hypersurface of constant energy $E$ in the $M$-sampling ensemble.
For the sampling ensemble, the $M$-dependence of density of states, entropy, and temperature implies that sample size should be treated as a independent state variable. In the large-$M$ limit, the sampling ensemble corresponds to Hill’s nanothermodynamics with $\mathcal{N}$-copies, as realized through the replica trick [41, 42, 43, 44, 45, 21]. Accordingly, thermodynamic quantities of entropy and temperature coincide with their counterparts in nanothermodynamics.
Because $\hat{\beta}_{B,M}$ is the UMVUE of $\beta$ in $M$ samplings, its variance is the infimum of among variances of unbiased estimators of $\beta$, i.e.,
$\sqrt{M}\Delta\hat{\beta}_{M}\Delta E\geq\sqrt{M}\Delta\hat{\beta}_{B,M}\Delta E\geq 1,$ (24)
where $\Delta E=\Delta E_{\bm{i}}/\sqrt{M}$ denotes the standard deviation of energy in the single sampling. Eq. (24) gives the achievable ETU in $M$ samplings. For large but finite $M$, the optimal estimation of temperature leads to a refined ETU following the similar procedure in Appendix E,
$\begin{split}\sqrt{M}\Delta\hat{\beta}_{M}\Delta E&\geq\sqrt{M}\Delta\hat{\beta}_{B,M}\Delta E\\ &=1+\frac{1}{4M}(\mu_{3})^{2}+O\left((MN)^{-2}\right).\end{split}$ (25)
Similarly, because $\hat{T}_{G,M}$ is the UMVUE of $T$ in $M$ samplings, and we obtain
$\frac{\sqrt{M}\Delta\hat{T}_{M}\Delta E}{T^{2}}\geq\frac{\sqrt{M}\Delta\hat{T}_{G,M}\Delta E}{T^{2}}\geq 1,$ (26)
and
$\begin{split}\frac{\sqrt{M}\Delta\hat{T}_{M}\Delta E}{T^{2}}&\geq\frac{\sqrt{M}\Delta\hat{T}_{G,M}\Delta E}{T^{2}}\\ &=1+\frac{T^{2}}{4MC^{3}}\left(\frac{\text{d}C}{\text{d}T}\right)^{2}+O\left((MN)^{-2}\right)\end{split}$ (27)
in the large $M$ limit. As $M\to\infty$, the Gaussian-type temperature fluctuation and the Cramér-Rao bound are recovered due to CLT. In Table 1, we also list the relevant quantities in the estimation of temperature for the single sampling and repeated sampling.
Table 1: Relevant quantities in estimation of temperature for the single sampling and the repeated sampling.
The single sampling The repeated sampling
Partition function $\ln Z=\ln\sigma(E^{*})-\beta E^{*}+\ln\sqrt{\left.2\pi\left(-\frac{\partial\hat{\beta}_{B}}{\partial E}\right)^{-1}\right|_{E=E^{*}}}+O(N^{-1})$ $M\ln Z=\ln\sigma_{M}(E_{\bm{i}}^{*})-\beta E_{\bm{i}}^{*}+\ln\sqrt{\left.2\pi\left(-\frac{\partial\hat{\beta}_{B,M}}{\partial E}\right)^{-1}\right|_{E_{t}=E_{t}^{*}}}+O\left((MN)^{-1}\right)$
Average energy $\left\langle E\right\rangle=E^{*}+\left.\frac{1}{2}\frac{\partial^{2}\hat{\beta}_{B}}{\partial E^{2}}\right|_{E=E^{*}}\left(\frac{\partial E^{*}}{\partial\beta}\right)^{2}+O(N^{-1})$ $M\left\langle E\right\rangle=E_{\bm{i}}^{*}+\left.\frac{1}{2}\frac{\partial^{2}\hat{\beta}_{B}}{\partial E_{\bm{i}}^{2}}\right|_{E_{\bm{i}}=E_{\bm{i}}^{*}}\left(\frac{\partial E_{\bm{i}}^{*}}{\partial\beta}\right)^{2}+O\left((MN)^{-1}\right)$
The most probable energy $E^{*}=\left\langle E\right\rangle-T+\frac{T^{2}}{2}\frac{\partial\ln C}{\partial T}+O(N^{-1})$ $E_{\bm{i}}^{*}=M\left\langle E\right\rangle-T+\frac{T^{2}}{2}\frac{\partial\ln C}{\partial T}+O\left((MN)^{-1}\right)$
Relation between canonical ensemble entropy and microcanonical ensemble entropy $\ln\sigma(\left\langle E\right\rangle)=S-\ln\sqrt{2\pi\left(-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)}+O(N^{-1})$ $\ln\sigma_{M}(M\left\langle E\right\rangle)=MS-\ln\sqrt{2\pi M\left(-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)}+O\left((MN)^{-1}\right)$
Relation between the UMVUE and the maximum likelihood estimator of the inverse temperature $\hat{\beta}_{B}(E)=\hat{\beta}_{L}(E)+\frac{1}{2}\frac{d^{2}_{E}\hat{\beta}_{L}(E)}{d_{E}\hat{\beta}_{L}(E)}+O(N^{-2})$ $\hat{\beta}_{B,M}(E)=\hat{\beta}_{L}(E)+\frac{1}{2M}\frac{d^{2}_{E}\hat{\beta}_{L}(E)}{d_{E}\hat{\beta}_{L}(E)}+O\left((MN)^{-2}\right)$
Refined ETU $\Delta\hat{\beta}_{B}\Delta E=1+\frac{1}{4}(\mu_{3})^{2}+O(N^{-2})$ $\sqrt{M}\Delta\hat{\beta}_{B,M}\Delta E=1+\frac{1}{4M}(\mu_{3})^{2}+O\left((MN)^{-2}\right)$
Refined $\Delta T\Delta E/T^{2}$ $\frac{\Delta\hat{T}_{G}\Delta E}{T^{2}}=1+\frac{T^{2}}{4C^{3}}(\frac{\text{d}C}{\text{d}T})^{2}+O(N^{-2})$ $\sqrt{M}\frac{\Delta\hat{T}_{G,M}\Delta E}{T^{2}}=1+\frac{T^{2}}{4MC^{3}}(\frac{\text{d}C}{\text{d}T})^{2}+O\left((MN)^{-2}\right)$
In FIG.5, we show the probability distributions of the estimators $\hat{T}_{G,M}$ and $\hat{T}_{B,M}=(\partial\ln[\sigma_{M}(E)\epsilon]/\partial E)^{-1}$ in finite-sized system for different $M$. For small $M$, they both follow a non-Gaussian distribution. The non-Gaussian probability distribution of $\hat{T}_{G,M}$ exactly corresponds to the steady state described in Ref. [49], which is a quasi-equilibrium state [50, 52, 53]. As $M$ increases, a Gaussian distribution naturally emerges. Meanwhile, regardless of the value of $M$, the optimal estimation of temperature is always unbiased ( $\left\langle\hat{T}_{G,M}\right\rangle=1$, while $\left\langle\hat{T}_{B,M}\right\rangle\neq 1$). Moreover, for $M=10$, the three estimators $\hat{T}_{G,1},\hat{T}_{G,5},\hat{T}_{G,10}$ are all unbiased. And the variance of $\hat{T}_{G,10}$ are minimum (its distribution is narrowest). Hence, our work provides a foundation for further experimental studies on the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature in finite-sized systems.
Figure 5: Probability distributions $P$ of the estimators $\hat{T}_{G,M}$ and $\hat{T}_{B,M}$ for different $M$, where the vertical lines indicate the mean values of the corresponding distributions. Here, $\sigma_{M}(E)=E^{\frac{3NM}{2}-1}$ is the density of states of the classical ideal gas [67]. We set temperature $T=1$ and particle number $N=6$.
V Conclusion and Outlook
In this work, we develop a systematic mathematical framework for measuring temperature fluctuations in finite-sized systems using the parameter estimation theory. Using this framework, we obtained the optimal estimation of temperature. It reveals the one-to-one link between different formulas of entropy and optimal estimation of different parameters. The UMVUE of temperature provides us with an achievable ETU, which is tighter than the usual one. In the large-$N$ limit, the optimal estimation of temperature leads to a refined ETU, which is related to the skewness of the canonical ensemble. Finally, as the sample size grows, the non-Gaussian distribution of temperature gradually approaches a Gaussian distribution due to the CLT, while the optimal estimation of temperature remains both unbiased and efficient—regardless of system size. Also, we provide a repeated sampling interpretation for the temperature and entropy in Hill’s nanothermodynamics.
Our work provides a theoretical foundation for further exploring the use of many-body systems to construct high-precision thermometers. The achievable bound for the ETU, the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature enable experimental testing in finite-sized systems, such as neutral atom arrays [74, 75] and biochemical oscillators [76, 77]. For thermometry, the optimal estimation of temperature may improve the validity of the measurement and the sampling efficiency in many-body systems.
Acknowledgements.
This work was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2024ZD0301000), the Science Challenge Project (No. TZ2025017), the National Natural Science Foundation of China (Grant No. 12405046), and the Science Foundation of Zhejiang Sci-Tech University (Grants No. 23062088-Y and No. 23062181-Y).
Appendix A Unbiased estimator of $\beta^{k}$
We prove that $\hat{\beta}^{(k)}$ is an unbiased estimator of $\beta^{k}$. Let us consider the case $k>0$. For $k=1$, $\text{d}_{E_{i}}\sigma(E_{i})=\text{d}\sigma(E_{i})/\text{d}E_{i}$, we have [78]
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\int_{-\infty}^{\infty}\text{d}E\delta(E-E_{i})\\ &=\frac{1}{Z}\int_{-\infty}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\sigma(E)}\frac{\text{d}\sigma(E)}{\text{d}E}\sum_{i\in I}\delta(E-E_{i})d_{i}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}{\text{e}^{-\beta E}}\text{d}\sigma(E)\\ &=\left.\frac{\text{e}^{-\beta E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\beta\text{e}^{-\beta E}\text{d}E\\ &=\beta,\end{split}$ (28)
where $\sigma(E)=\sum_{i}\delta(E-E_{i})d_{i}$ denotes the density of state, and we require that $\sigma(E_{\text{min}})=0$. In the derivation, the lower limit of the integral changes from negative infinity to $E_{\text{min}}$ due to $\sigma(E)=0$ for $E<0$. Eq. (7) was noted by Gibbs [79], but its connection to estimation theory was not explored.
Suppose that for $k=n$, $n$ is a positive integer, we have $\langle\hat{\beta}^{(n)}\rangle=\beta^{n}$. Then, for $k=n+1$, we have $\text{d}^{n+1}_{E_{i}}\sigma(E_{i})=\text{d}\left[\text{d}^{n}_{E_{i}}\sigma(E_{i})\right]/\text{d}E_{i}$ and
$\begin{split}\langle\hat{\beta}^{(n+1)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\left[\text{d}^{(n)}_{E_{i}}\sigma(E_{i})\right]}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{e}^{-\beta E}\text{d}\left[\text{d}^{n}_{E}\sigma(E)\right]\\ &=\left.\frac{\text{e}^{-\beta E}\text{d}^{n}_{E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{d}\text{e}^{-\beta E}\\ &=\beta\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{e}^{-\beta E}\text{d}E\\ &=\beta^{n+1},\end{split}$ (29)
where we require that $\left.\text{d}^{l}_{E}\sigma(E)\right|_{E=E_{\text{min}}}=0$ for integer $l\in\left[0,n-1\right]$. Therefore, by using the method of mathematical induction, we prove that $\hat{\beta}^{(k)}=\text{d}^{k}_{E_{i}}\sigma(E_{i})/\sigma(E_{i})$ is an unbiased estimate of $\beta^{k}$ when $k>0$.
Next, let us consider the case $k<0$. For k=1k=-1, $\text{d}_{E_{i}}^{-1}\sigma(E_{i})=\int_{E_{\text{min}}}^{E_{i}}\text{d}E^{\prime}\sigma(E^{\prime})$, we have
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\frac{1}{Z}\sum_{i\in I}\frac{\text{d}_{E_{i}}^{-1}\sigma(E_{i})}{\sigma(E_{i})}d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{-1}{Z\beta}\int_{E_{\text{min}}}^{\infty}\left(\text{d}_{E}^{-1}\sigma(E)\right)\text{d}\text{e}^{-\beta E}\\ =&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{\infty}_{E_{\text{min}}}\\ &+\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{1}{\beta}\int_{E_{\text{min}}}^{\infty}\sigma(E)\frac{\text{e}^{-\beta E}}{Z}\text{d}E\\ =&\beta^{-1}.\end{split}$ (30)
Follow the similar procedure, we prove that $\hat{\beta}^{(k)}=\text{d}_{E_{i}}^{k}\sigma(E_{i})/\sigma(E_{i})$ is an unbiased estimate of $\beta^{k}$ when $k<0$.
Appendix B The density of states of an analogous low-temperature ideal Fermi gas
The internal energy of an low-temperature ideal Fermi gas is given by (see Section 8.1 of Ref. [69])
$E=\frac{3}{5}N\varepsilon_{F}+\frac{\pi^{2}}{4}N\varepsilon_{F}\left(\frac{T}{\varepsilon_{F}}\right)^{2}=E_{\text{min}}+aNT^{2},$ (31)
and the entropy is given by
$S=\frac{\pi^{2}TN}{2\varepsilon_{F}}=bNT,$ (32)
where $\varepsilon_{F}$ is generally referred to as the Fermi energy of the system, $a=\pi^{2}/4\varepsilon_{F}$ and $b=\pi^{2}/2\varepsilon_{F}$. Then, $T$ is obtained from Eq. (31):
$T=\sqrt{\frac{E-E_{\text{min}}}{aN}}.$ (33)
Using the Boltzmann relation $S=\ln\sigma(E)$ and Eq. (33), the density of states $\sigma(E)$ can be expressed as:
$\sigma(E)=\text{e}^{S}=\text{e}^{bNT}=\text{e}^{\frac{b}{\sqrt{a}}\sqrt{N(E-E_{\text{min}})}}.$ (34)
For simplicity, we set $b/\sqrt{a}=1$ when plotting FIG. 1. However, the qualitative results are unaffected regardless of the value of $b/\sqrt{a}$. The above approach provides an approximation for the density of states of a low-temperature ideal Fermi gas in the limit of $N\to\infty$. For finite $N$, a more detailed discussion of the density of states can be found in Refs. [80, 81].
Appendix C Relative biases of $\beta$ and $T$ in a two-level system
Firstly, for systems with finite $E_{\text{max}}$, Eq. (29) is modified to
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\left.\frac{\sigma(E)\text{e}^{-\beta E}}{Z}\right|^{E_{\text{max}}}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{\sigma(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\sigma(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{Z}+\beta.\end{split}$ (35)
In an $N$ non-interacting two-level system, both the upper bound $E_{\text{max}}=N\varepsilon/2$ and lower bound $E_{\text{min}}=-N\varepsilon/2$ bounds of system’s energy are finite. Here [69],
$\sigma(E)=\frac{\Gamma(N+1)}{\varepsilon\Gamma\left(\frac{N}{2}-\frac{E}{\varepsilon}+1\right)\Gamma\left(\frac{N}{2}+\frac{E}{\varepsilon}+1\right)}$ (36)
and $Z=(\text{e}^{\beta\varepsilon/2}+\text{e}^{-\beta\varepsilon/2})^{N}$. Then, Eq. (35) is
$\langle\hat{\beta}^{(1)}\rangle=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta,$ (37)
And the relative bias of $\beta$ is
$\begin{split}\Phi_{\beta}&=\left|\frac{\langle\hat{\beta}^{(1)}\rangle-\beta}{\beta}\right|\\ &=\left|\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\beta\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}\right|\\ &\approx\frac{1}{\beta\varepsilon(1+\text{e}^{-|\beta|\varepsilon})^{N}}\\ &=\frac{1}{\beta\varepsilon}\text{e}^{-N(\ln(1+\text{e}^{-|\beta|\varepsilon}))}\\ &\sim\text{e}^{-\alpha_{1}N},\end{split}$ (38)
where the approximation used applies $\text{e}^{-N|\beta|\varepsilon}\ll 1$ with large $N$. In FIG. 6a, $\alpha_{1}$ decreases exponentially with increasing inverse temperature. For a particular $N$, $\Phi_{\beta}$ is non-negligible, which indicates that the unbiased estimator $\hat{\beta}^{(1)}$ remains unbiased except at extremely low temperatures, as given in Eq. (6).
Figure 6: Decay rate varies as NN when $E_{\text{max}}$ is finite. Here, the density of states of an $N$ non-interacting two-level system is Eq. (36) and we set the spacing of energy levels $\varepsilon=1$.
Secondly, for systems with finite $E_{\text{max}}$, Eq. (30) is modified to
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{E_{\text{max}}}_{E_{\text{min}}}\\ &+\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{\Omega(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\Omega(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{\beta Z}+\beta^{-1}.\end{split}$ (39)
In an $N$ non-interacting two-level system, we have $\Omega(E_{\text{min}})=\int_{E_{\text{min}}}^{E_{\text{min}}}\sigma(E)\text{d}E=0$ and $\Omega(E_{\text{max}})=\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}E=Z|_{\beta=0}$. Then, with the help of Eq. (36), Eq. (39) is
$\langle\hat{\beta}^{(-1)}\rangle=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\beta\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta^{-1},$ (40)
thus, the relative bias of $T$ is
$\begin{split}\Phi_{T}&=\left|\frac{\langle\hat{\beta}^{(-1)}\rangle-T}{T}\right|\\ &=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}})^{N}}\\ &=\left(\frac{2\text{e}^{\frac{-\beta\varepsilon}{2}}}{\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}}\right)^{N}\\ &=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\cosh(\frac{\beta\varepsilon}{2})^{N}}\\ &=\text{e}^{-N(\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2}))}\\ &\sim\text{e}^{-\alpha_{2}N}.\end{split}$ (41)
In FIG. 6b, when $T<0$, $\alpha_{2}$ becomes negative, causing $\Phi_{T}$ to increase exponentially, and the unbiased estimator of $\hat{\beta}^{(-1)}$ becomes invalid. When $T>0$, $\alpha_{2}$ is positive, and the unbiased estimator of $\hat{\beta}^{(-1)}$ is valid, except at extremely high temperatures where $\alpha_{2}$ approaches zero.
Appendix D Unbiased estimator of $\beta^{\alpha}$ for $\alpha\in\mathbb{R}$
In this appendix, let us consider the unbiased estimation of the non-integer powers of temperature, i.e., $\langle\hat{\beta}^{(\alpha)}\rangle=\beta^{\alpha}$ for $\alpha\in R$. The expression of $\hat{\beta}^{(\alpha)}$ is still given by Eq. (6) except that we extend the definition of $\text{d}^{\alpha}_{E}\sigma(E)$ in Eq. (7) in different cases. When $\alpha<0$ , the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\sigma(u)\left(E-t\right)^{-\alpha-1}\text{d}u,$ (42)
using the Riemann-Liouville integral [82]. And we have
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (43)
When $\alpha>0$ and $n-1<\alpha<n$ for some integer $n$, the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{n-\alpha-1}\text{d}^{n}_{u}\sigma(u)\text{d}u,$ (44)
using the Caputo fractional derivative [83] under the condition $\left.\text{d}_{E}^{l}\sigma(E)\right|_{E=0}=0$ for integer $l\in\left[0,n-1\right]$. And we have
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle=&\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(n-\alpha)}\\ &\int_{E_{\text{min}}}^{E}\text{d}t\text{d}_{t}^{n}\sigma(t)\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\\ &\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{\beta^{\alpha-n}}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ =&\frac{1}{Z}\beta^{\alpha-n}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\\ =&\beta^{\alpha}.\end{split}$ (45)
The last equlity is obtained by using Eq. (29).
When $\alpha>0$ and the condition $\left.\text{d}_{E}^{l}\sigma(E)\right|_{E=0}=0$ is not satisfied for every $l\in\left[0,n-1\right]$, we only consider the case $\alpha\in\left[0,1\right)$. Here, the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(1-\alpha)}\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{-\alpha}\sigma(u)\text{d}u,$ (46)
using the Riemann-Liouville fractional derivative [82]. And we have
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(1-\alpha)}\\ &\quad\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\left.\frac{1}{\Gamma(1-\alpha)}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\right|^{\infty}_{0}\\ &\quad-\frac{1}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (47)
We would like to emphasize that in the limit $\alpha\to 1^{-}$, $\lim_{\alpha\to 1^{-}}\langle\hat{\beta}^{(\alpha)}\rangle=\beta$, whereas $\langle\hat{\beta}^{(1)}\rangle=\beta+\sigma(E_{\text{min}})/Z$. This means the interchange of the limit $\alpha\to 1^{-}$ and the integral $\left\langle\cdot\right\rangle$ is not allowed.
Appendix E The approximate expression of $\hat{\beta}_{B}$ and $\hat{T}_{G}$ in large-$N$ limit
Using the Laplace approximation [72, 73] or the Darwin–Fowler method in Sec. 9.1 of [67], $\ln Z$ can be approximated as [84]
$\begin{split}\ln Z=&\ln\sigma(E^{*})+\ln\sqrt{\left.2\pi\left(-\frac{\partial\hat{\beta}_{B}}{\partial E}\right)^{-1}\right|_{E=E^{*}}}\\ &-\beta E^{*}+O\left(N^{-1}\right),\end{split}$ (48)
where $\left.\partial\ln\sigma(E)/\partial E\right|_{E=E^{*}}=\beta$, i.e. $\hat{\beta}_{B}(E^{*})=\beta$. Here, $E^{*}$ denotes the most probable energy in the canonical ensemble. Then, we obtain the expression average energy $\left\langle E\right\rangle=-\partial\ln Z/\partial\beta$,
$\left\langle E\right\rangle=E^{*}+\left.\frac{1}{2}\frac{\partial^{2}\hat{\beta}_{B}}{\partial E^{2}}\right|_{E=E^{*}}\left(\frac{\partial E^{*}}{\partial\beta}\right)^{2}+O(N^{-1}).$ (49)
Performing the inverse map from $\left\langle E\right\rangle$ to $E^{*}$ shows the relation between the most probable energy $E^{*}$ and the average energy $\left\langle E\right\rangle$,
$E^{*}=\left\langle E\right\rangle-T+\frac{T^{2}}{2}\frac{\partial\ln C}{\partial T}+O(N^{-1}),$ (50)
where $C=\partial\left\langle E\right\rangle/\partial T$ denotes the heat capacity. Substituting Eq. (50) into Eq. (48), we have
$\ln\sigma(\left\langle E\right\rangle)=S-\ln\sqrt{2\pi\left(-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)}+O(N^{-1}),$ (51)
where $S=\ln Z-\beta\partial\ln Z/\partial\beta$ is the thermodynamics entropy in a canonical ensemble. Eq. (51) shows the nonequivalence between micro-canonical ensemble and canonical ensemble for finite-$N$ systems, i.e., the relation between the Boltzmann entropy $S_{B}=\ln[\sigma(E)\epsilon]$ and the thermodynamic entropy $S(E)$. If we replace $\left\langle E\right\rangle$ with a sample of system’s energy $E$, the function $\hat{\beta}_{L}(E)$ (given by $\left.\frac{\partial\ln Z}{\partial\beta}\right|_{\beta=\hat{\beta}_{L}(E)}=E$) will be a statistic of $\beta$ [85]. Taking the derivative of Eq. (51) with respect to $E$ on both sides, we have
$\hat{\beta}_{B}(E)=\hat{\beta}_{L}(E)+\frac{1}{2}\frac{\text{d}^{2}_{E}\hat{\beta}_{L}(E)}{\text{d}_{E}\hat{\beta}_{L}(E)}+O(N^{-2}).$ (52)
This is the relation between the UMVUE of $\beta$ and the first-order moment $\hat{T}_{L}(E)=\hat{\beta}_{L}(E)^{-1}$ (also the maximum likelihood) estimator of $\beta$.
Performing Taylor expansion of Eq. (52) at $\langle E\rangle$ with $\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)=\beta$, and considering $\delta E=E-\langle E\rangle$ as a small quantity, we have
$\begin{split}\hat{\beta}_{B}-\beta=&\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\delta E+\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{2}}{2}\\ &+\text{d}_{\left\langle E\right\rangle}^{3}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{3}}{6}+\frac{1}{2}\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}\\ &+\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}\right)\frac{\delta E}{2}+O(N^{-2}).\end{split}$ (53)
Using Eq. (53), we obtain the variance of $\hat{\beta}_{B}$, $\Delta\beta^{2}_{B}=\langle(\hat{\beta}_{B}-\beta)^{2}\rangle$, which can be expanded as
$\begin{split}\Delta\beta^{2}_{B}=&\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}\left\langle(\delta E)^{2}\right\rangle+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left\langle(\delta E)^{3}\right\rangle\\ &+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)\left\langle(\delta E)^{2}\right\rangle\\ &+\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{2}\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\left\langle(\delta E)^{2}\right\rangle+\frac{1}{4}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)^{2}\\ &+\frac{\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)}{3}\left\langle(\delta E)^{4}\right\rangle+\frac{\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}}{4}\left\langle(\delta E)^{4}\right\rangle\\ &+O(N^{-3}).\end{split}$ (54)
In the derivation, we have used $\delta E\sim N^{1/2}$ and $\text{d}_{\left\langle E\right\rangle}\sim N^{-1}$.
To simplify Eq. (54), we introduce the higher order derivatives of the inverse function [86],
$\begin{split}\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-1}\\ \text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=-\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-3}\\ \text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-5}\left[3\left(\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\right)^{2}-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\frac{\partial^{3}\left\langle E\right\rangle}{\partial\beta^{3}}\right]\end{split}$ (55)
Then, substituting Eq. (55) into Eq. (54), we obtain the variance of $\hat{\beta}_{B}$,
$\Delta\hat{\beta}_{B}^{2}=\frac{1}{(\Delta E)^{2}}+\frac{1}{2}\frac{\left\langle(\delta E)^{3}\right\rangle^{2}}{(\Delta E)^{8}}+O(N^{-3}).$ (56)
Next, we consider the approximate expression of $S_{G}=\ln\Omega$. Using the relation $\partial\Omega(E)/\partial E=\sigma(E)$, we have,
$S_{G}(E)=S_{B}(E)+\ln\hat{T}_{G}(E)+O(N^{-1}),$ (57)
Eq. (57) is consistent with the exact relation $\hat{T}_{G}=\hat{T}_{B}(1-C_{G}^{-1})$ , where $C_{G}=\partial E/\partial\hat{T}_{G}$ [1]. Taking the derivative of Eq. (57) with respect to $E$ on both sides and using Eq. (52), we have
$\hat{T}_{G}(E)=\hat{T}_{L}(E)-\frac{\hat{T}_{L}(E)^{2}}{2}\frac{\text{d}^{2}_{E}\hat{T}_{L}(E)}{\text{d}_{E}\hat{T}_{L}(E)}+O(N^{-2}),$ (58)
Following the similar procedure, we obtain the variance of $\hat{T}_{G}$ using Eqs. (55) (58) and $\hat{T}_{L}\left(\left\langle E\right\rangle\right)=T$,
$\Delta T^{2}_{G}=\frac{T^{4}}{(\Delta E)^{2}}+\frac{T^{12}(\text{d}^{2}_{T}\left\langle E\right\rangle)^{2}}{2(\Delta E)^{8}}+O(N^{-3}).$ (59)
Appendix F Estimation of temperature for the classical ideal gas
In the classical ideal gas, the density of states is $\sigma(E)=E^{3N/2-1}$ [67]. We obtain the partition function of the system
$Z=\int_{0}^{\infty}\sigma(E)\text{e}^{-\beta E}=\beta^{-\frac{3N}{2}}\Gamma\left(\frac{3N}{2}\right),$ (60)
the most probable energy
$E^{*}=\frac{3N-2}{2\beta},$ (61)
the Boltzmann temperature
$\hat{\beta}_{B}=\frac{3N-2}{2E},$ (62)
and the Gibbs temperature
$\hat{T}_{G}=\frac{2E}{3N}.$ (63)
Using Eq. (60), we obtain the average energy
$\left\langle E\right\rangle=-\frac{\partial\ln Z}{\partial\beta}=\frac{3N}{2\beta}.$ (64)
the heat capacity
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2}$ (65)
and the first moment estimator of the temperature (inverse temperature)
$\hat{T}_{L}(E)=\frac{2E}{3N},\quad\hat{\beta}_{L}(E)=\frac{3N}{2E}.$ (66)
Substitute the above expressions into both sides of Eqs. (11), (12) and (15). We have
$\begin{split}&\left(\frac{3N}{2}-1\right)\left(\ln\frac{3N}{2}-\ln\beta\right)\\ =&\left(1-\frac{3N}{2}\right)\ln\beta+\ln\Gamma\left(\frac{3N}{2}\right)-\frac{1}{2}\ln 2\pi\\ &-\frac{1}{2}\ln\left(\frac{3N}{2}\right)+\left(\frac{3N}{2}\right)+O(N^{-1}),\end{split}$ (67)
where we have used the String approximation [87] in the derivation.
Moreover, the left-hand side (LHS) and the right-hand side (RHS) of Eqs. (12) and (15) are exactly equal, because our model is sufficiently simple. In general, they are approximately valid.
The probability distribution of the system’s energy is a gamma distribution
$P(E)=\frac{\beta^{\frac{3N}{2}}E^{\frac{3N}{2}-1}\text{e}^{-\beta E}}{\Gamma\left(\frac{3N}{2}\right)}.$ (68)
Using Eq. (68), we obtain the standard deviation of Boltzmann temperature
$\Delta\hat{\beta}_{B}=\beta\sqrt{\frac{2}{3N-4}},$ (69)
the standard deviation of energy
$\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (70)
and the skwness of this disribution
$\mu_{3}=\sqrt{\frac{8}{3N}}.$ (71)
Thus, the result of the achievable ETU (Eq. (13)) using Eqs. (69) (70) is:
$\begin{split}\Delta\hat{\beta}\Delta E&\geq\Delta\hat{\beta}_{B}\Delta E=\sqrt{\frac{3N}{3N-4}}>1.\end{split}$ (72)
And the refined ETU (Eq. (14)) is expressed as
$\begin{split}\sqrt{\frac{3N}{3N-4}}&=1+\frac{2}{3N}+O(N^{-2}).\end{split}$ (73)
Following the similar procedure, we obtain
$\Delta\hat{T}_{G}=\beta^{-1}\sqrt{\frac{2}{3N}},\quad\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (74)
and
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2}.$ (75)
Then, Eqs. (16) and (17) are given straightforwardly.